COMPUTATIONAL APPROACHES TO THE DETERMINATION OF THE MOLECULAR GEOMETRY OF ACROLEIN IN ITS T_{1}(n, π^{∗}) STATE

\underline{MICHAEL O. MCANALLY}, NIKOLAUS C. HLAVACEK, and STEPHEN DRUCKER, Department of Chemistry, University of Wisconsin-Eau Claire, Eau Claire, WI 54702.

The spectroscopically derived inertial constants for acrolein (propenal) in its T_{1}(n, π^{∗}) state^{a} were used to test predictions from a variety of computational methods. One focus was on multiconfigurational methods, such as CASSCF and CASPT2, that are applicable to excited states. We also examined excited-state methods that utilize single reference configurations, including EOM-EE-CCSD and TD-PBE0. Finally, we applied unrestricted ground-state techniques, such as UCCSD(T) and the more economical UPBE0 method, to the T_{1}(n, π^{∗}) excited state under the constraint of C_{s} symmetry. The unrestricted ground-state methods are applicable because at a planar geometry, the T_{1}(n, π^{∗}) state of acrolein is the lowest-energy state of its spin multiplicity. Each of the above methods was used with a triple zeta quality basis set to optimize the T_{1}(n, π^{∗}) geometry. This procedure resulted in the following sets of inertial constants:

Inertial constants (cm^{−1}) of acrolein in its T_{1}(n, π^{∗}) state

Method  A  B  C  Method  A  B  C
CASPT2(6,5)^{b}  1.667 0.1491 0.1368 UCCSD(T)^{b}  1.668 0.1480 0.1360
CASSCF(6,5)  1.667 0.1491 0.1369 UPBE0  1.699 0.1487 0.1367
EOM-EE-CCSD 1.675 0.1507 0.1383
TD-PBE0  1.719 0.1493 0.1374 Experiment^{a} 1.662 0.1485 0.1363

The two multiconfigurational methods produce the same inertial constants, and those constants agree closely with experiment. However the sets of computed bond lengths differ significantly for the two methods. In the CASSCF calculation, the lengthening of the C=O and C=C bonds and the shortening of the C—C bond are more pronounced than in CASPT2.

aPrevious talk

bO. S. Bokareva et al., Int. J. Quant. Chem. 108, 2719 (2008).